# Solvability of fractional semilinear heat equations with distributional forcing terms in Lorentz spaces

### v1

@@ 3.2 Contraction mapping argument
Assume that $\mu$ satisfies $I[\mu]\in B_{M/2}$ and $p$ satisfies $p>\gamma$ and $p\geqslant\frac{N(\gamma-1)}{\theta}$. Then, the operator $u\mapsto\mathcal{K}[u]$ is a map from $B_{M}$ to $B_{M}$.
In the following proof, we divide the case into $p>\frac{N(\gamma-1)}{\theta}$ and $p=\frac{N(\gamma-1)}{\theta}$. While they can be discussed jointly with minor adjustments, we separate them here for simplicity.
@@
$p>\frac{N(\gamma-1)}{\theta}$ case. Take arbitrary $u\in B_{M}$. By utilizing Proposition 2.14 and Lemma 2.1 (2), we have
$\left\|J[u](\cdot,t)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\leqslant C\int_{0}^{t}\left\{1+(t-\tau)^{-\frac{N(\gamma-1)}{\theta p}}\right\}\left\||u|^{\gamma-1}u(\cdot,\tau)\mathrel{}\middle|\mathrel{}L^{p/\gamma,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\leqslant Ct^{1-\frac{N(\gamma-1)}{\theta p}}\left\|u\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma}\leqslant CT^{1-\frac{N(\gamma-1)}{\theta p}}M^{\gamma}\leqslant M/2$
for all $0<t<T\leqslant 1$. This gives $\mathcal{K}[u]\in B_{M}$.
$p=\frac{N(\gamma-1)}{\theta}$ case. Pick arbitrary $u\in B_{M}$. We take $\rho>0$ and $k(\rho)>0$ such that
$1<p/\gamma<\rho<p=\frac{N(\gamma-1)}{\theta}<k(\rho)\quad\text{and}\quad\frac{1}{p}=\frac{1/2}{\rho}+\frac{1/2}{k(\rho)}.$ (3.3)
@@
$\eta=\eta(\lambda):=\lambda^{b},\quad b:=\frac{\theta p\rho}{2(\theta p\rho+Np-N\gamma\rho)}$ (3.4)
@@
$J[u](x,t)$ $=\int_{0}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\,d\tau$
$=\left(\int_{0}^{\eta}+\int_{\eta}^{\infty}\right)\chi_{\left\{0<\tau<t\right\}}(\tau)S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\,d\tau$
$=:J_{\eta}^{0}[u](x,t)+J_{\eta}^{1}[u](x,t).$
By Proposition 2.14, (3.3) and (3.4), we have
$\left\|J_{\eta}^{0}[u](x,t)\mid L^{\rho,\infty}_{\mathrm{ul}}\right\|$ $\leqslant\int_{0}^{\eta}\chi_{\left\{0<\tau<t\right\}}(\tau)\left\|S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\mid L^{\rho,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\leqslant C\int_{0}^{\eta}\left(1+\tau^{\frac{N}{\theta}\left(\frac{1}{\rho}-\frac{\gamma}{p}\right)}\right)\chi_{\left\{0<\tau<t\right\}}(\tau)\left\||u|^{\gamma-1}u(\cdot,t-\tau)\mid L^{p/\gamma,\infty}_{\mathrm{ul}}\right\|d\tau$
$=C\eta^{\frac{N}{\theta}\left(\frac{1}{\rho}-\frac{\gamma}{p}\right)+1}\left\|u\mid L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma}\leqslant C\eta^{\frac{N}{\theta}\left(\frac{1}{\rho}-\frac{\gamma}{p}\right)+1}M^{\gamma}$
$=C\lambda^{1/2}M^{\gamma}$
and
$\left\|J_{\eta}^{1}[u](x,t)\mid L^{k(\rho),\infty}_{\mathrm{ul}}\right\|$ $\leqslant\int_{\eta}^{\infty}\chi_{\left\{0<\tau<t\right\}}(\tau)\left\|S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\mid L^{k(\rho),\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\leqslant C\int_{\eta}^{\infty}\tau^{\frac{N}{\theta}\left(\frac{1}{k(\rho)}-\frac{\gamma}{p}\right)}\left\||u|^{\gamma-1}u(\cdot,t-\tau)\mid L^{p/\gamma,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$=C\eta^{\frac{N}{\theta}\left(\frac{1}{k(\rho)}-\frac{\gamma}{p}\right)+1}\left\|u\mid L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma}\leqslant C\eta^{\frac{N}{\theta}\left(\frac{1}{k(\rho)}-\frac{\gamma}{p}\right)+1}M^{\gamma}$
$=C\lambda^{-1/2}M^{\gamma}.$
The use of Corollary 2.8 in the form $L^{p,\infty}_{\mathrm{ul}}\supset\left(L^{\rho,\infty}_{\mathrm{ul}},L^{k(\rho),\infty}_{\mathrm{ul}}\right)_{1/2,\infty}$ implies that
$\left\|J[u](\cdot,t)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\sup_{\lambda>0}\lambda^{-1/2}\inf_{J[\mu]=J_{0}+J_{1}}\left\{\left\|J_{0}\mathrel{}\middle|\mathrel{}L^{\rho,\infty}_{\mathrm{ul}}\right\|+\lambda\left\|J_{1}\mathrel{}\middle|\mathrel{}L^{k(\rho),\infty}_{\mathrm{ul}}\right\|\right\}$
$\leqslant C\sup_{\lambda>0}\left\{\lambda^{-1/2}\left\|J_{\eta(\lambda)}^{0}[u]\mathrel{}\middle|\mathrel{}L^{\rho,\infty}_{\mathrm{ul}}\right\|+\lambda^{1/2}\left\|J_{\eta(\lambda)}^{1}[u]\mathrel{}\middle|\mathrel{}L^{k(\rho),\infty}_{\mathrm{ul}}\right\|\right\}$
@@
$\sup_{0<t<T}\left\|J[u](\cdot,t)\mid L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant\frac{M}{2}$
@@
Assume that $p$ satisfies the same conditions as in Lemma 3.3. Then, the map $\mathcal{K}:B_{M}\to B_{M}$ is a contraction map.
@@
Firstly by using Lemma 2.1, we prepare the following estimate. For all $q>1$, $\sigma\geqslant 1$ and for all $f$, $g\in L^{q\sigma,\infty}_{\mathrm{ul}},$ we have
$\left\||f|^{q-1}f-|g|^{q-1}g\mathrel{}\middle|\mathrel{}L^{\sigma,\infty}_{\mathrm{ul}}\right\|$ $\leqslant q\left\|\max\{|f|;|g|\}^{q-1}|f-g|\mathrel{}\middle|\mathrel{}L^{\sigma,\infty}_{\mathrm{ul}}\right\|$
$\leqslant q\left\|\max\{|f|;|g|\}^{q-1}\mathrel{}\middle|\mathrel{}L^{\frac{q\sigma}{q-1},\infty}_{\mathrm{ul}}\right\|\left\|f-g\mid L^{q\sigma,\infty}_{\mathrm{ul}}\right\|$
$\leqslant q\left(\left\|f\mathrel{}\middle|\mathrel{}L^{q\sigma,\infty}_{\mathrm{ul}}\right\|^{q-1}+\left\|g\mathrel{}\middle|\mathrel{}L^{q\sigma,\infty}_{\mathrm{ul}}\right\|^{q-1}\right)\left\|f-g\mid L^{q\sigma,\infty}_{\mathrm{ul}}\right\|$. (3.6)
$p>\frac{N(\gamma-1)}{\theta}$ case. Take any $u,v\in B_{M}$. For all $0<t<T\leqslant 1$, it holds that
$\left\|\mathcal{K}[u](\cdot,t;\mu)-\mathcal{K}[v](\cdot,t;\mu)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|=\left\|J[u](\cdot,t)-J[v](\cdot,t)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\quad\leqslant C\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)-|v|^{\gamma-1}v(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\quad\leqslant C\int_{0}^{t}\left\{1+(t-\tau)^{-\frac{N(\gamma-1)}{\theta p}}\right\}\left\||u|^{\gamma-1}u(\cdot,\tau)-|v|^{\gamma-1}v(\cdot,\tau)\mathrel{}\middle|\mathrel{}L^{p/\gamma,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\quad\leqslant C\int_{0}^{t}(t-\tau)^{-\frac{N(\gamma-1)}{\theta p}}\left(\left\|u\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma-1}+\left\|v\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|^{\gamma-1}\right)\left\|u-v\mid L^{p,\infty}_{\mathrm{ul}}\right\|\,d\tau$
$\quad\leqslant Ct^{1-\frac{N(\gamma-1)}{\theta p}}M^{\gamma-1}\left\|u-v\mid X_{T}\right\|\leqslant\frac{1}{2}\left\|u-v\mid X_{T}\right\|$
through the use of (3.2) and Proposition 2.14.
$p=\frac{N(\gamma-1)}{\theta}$ case. Utilizing (3.2) and carrying out a similar procedure as in the proof of Lemma 3.3, the claim follows. ∎
@@
From Lemma 3.3, Lemma 3.4 and the contraction mapping theorem, we only have to show $\left\|I[\mu]\mathrel{}\middle|\mathrel{}X_{T}\right\|\leqslant M/2$. Assertion (1) is proved by Proposition 3.2. We next prove (3). This is established by following calculation. We use Proposition 2.9 and 2.13 to see that
$\left\|I[\mu](\cdot,t)\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{0}_{(p,\infty),1}\right\|\,d\tau$
$\leqslant C\int_{0}^{t}\left\{1+(t-\tau)^{s/\theta}\right\}\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|\,d\tau$
$\leqslant C\left(t+t^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|\leqslant C\left(T+T^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|$
@@
$\left\|I[\mu]\mathrel{}\middle|\mathrel{}X_{T}\right\|$ $\leqslant C\left(T+T^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s}_{(p,\infty),\infty}\right\|\leqslant\frac{M}{2}$ (3.7)
@@
$\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{\frac{Np}{N+p\varepsilon},\infty}_{\mathrm{ul}}\right\|<\delta.$
$\delta>0$ is determined later. By definition of $\limsup$, there exists an integer $m\in\mathbb{Z}$ such that $2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{\frac{Np}{N+p\varepsilon},\infty}_{\mathrm{ul}}\right\|<\delta$ holds for all $j\geqslant m$. Set $\mu_{1}:=\sum_{j=0}^{m}\Delta_{j}\mu$ and $\mu_{2}:=\mu-\mu_{1}$. Then for every $j\in\mathbb{N}$, we have the following relations.
$\Delta_{j}\mu_{1}$ $=\begin{cases}\Delta_{j}\mu&\text{for }j\leqslant m-1\\ \left(\Delta_{m-1}+\Delta_{m}\right)\Delta_{j}\mu&\text{for }j=m,m+1\\ 0&\text{for }j\geqslant m+2,\end{cases}$
$\Delta_{j}\mu_{2}$ $=\begin{cases}0&\text{for }j\leqslant m-1\\ \left(\Delta_{m+1}+\Delta_{m+2}\right)\Delta_{j}\mu&\text{for }j=m,m+1\\ \Delta_{j}\mu&\text{for }j\geqslant m+2.\end{cases}$
Thus we have $\left\|\mu_{2}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{\left(\frac{Np}{N+p\varepsilon},\infty\right),\infty}\right\|<\delta$ and
$\left\|\mu_{1}\mathrel{}\middle|\mathrel{}B^{\sigma}_{(p,\infty),\infty}\right\|$ $=\max_{0\leqslant j\leqslant m+1}2^{\sigma j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant\max_{0\leqslant j\leqslant m+1}2^{(\sigma+\theta)j}2^{-\theta j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant 2^{(\sigma+\theta)(m+1)}\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta}_{(p,\infty),\infty}\right\|\leqslant 2^{(\sigma+\theta)(m+1)}\left\|\mu\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{\left(\frac{Np}{N+p\varepsilon},\infty\right),\infty}\right\|$
for some $\sigma\in\left]-\theta,0\right[$. Now applying (3.7) and Proposition 3.1 to $\mu_{1}$ and $\mu_{2}$ respectively, we get
@@
$\leqslant C\left(T+T^{1+s/\theta}\right)\left\|\mu_{1}\mathrel{}\middle|\mathrel{}B^{\sigma}_{(p,\infty),\infty}\right\|+C_{\varepsilon}\left\|\mu_{2}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{(\frac{Np}{N+p\varepsilon},\infty),\infty}\right\|$
$\leqslant C\left(T+T^{1+s/\theta}\right)2^{\theta m}\left\|\mu\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{(\frac{Np}{N+p\varepsilon},\infty),\infty}\right\|+C_{\varepsilon}\delta.$
@@
Let $p$ satisfy the same conditions in Theorem 1.5. If the forcing term $\mu$ satisfies one of the conditions in Theorem 1.5, we find a number $s<0$ such that $\mu$ belongs to $B^{s}_{(p,\infty),\infty}\subset B^{s-N/p}_{\infty,\infty}$. Then the solution $u(x,t)$ constructed above satisfy the initial condition in the following sense: $u(x,t)\to 0$ in the weak-$\ast$ topology of the space $B^{s-N/p}_{\infty,\infty}$ as $t\to+0$. Take arbitrary $\psi\in B^{N/p-s}_{1,1}$ and calculate as follows:

### v2

@@ 3.2 Contraction mapping argument
Fix $0<T<1$. Assume that $\mu$ satisfies $I[\mu]\in B_{M/2}$ and $(p,q)$ satisfies $\gamma<q\leqslant p$ and $p\geqslant{N(\gamma-1)}/{\theta}$. Then the operator $u\mapsto\mathcal{K}[u;\mu]$ is a map from $B_{M}$ to $B_{M}$.
@@
Case $p>{N(\gamma-1)}/{\theta}$. Take arbitrary $u\in B_{M}$. By utilizing Proposition 2.16 and Lemma 2.2 (2), we have
$\left\|J[u](\cdot,t)\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ $\leqslant C\int_{0}^{t}\left\|S(t-\tau)\left[|u|^{\gamma-1}u(\cdot,\tau)\right]\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\,d\tau$
$\leqslant C\int_{0}^{t}\left\{1+(t-\tau)^{-\frac{N(\gamma-1)}{\theta p}}\right\}\left\||u|^{\gamma-1}u(\cdot,\tau)\mathrel{}\middle|\mathrel{}M^{p/\gamma}_{q/\gamma,\infty}\right\|\,d\tau$
$\leqslant CT^{1-\frac{N(\gamma-1)}{\theta p}}M^{\gamma}\leqslant M/2$
for all $0<t<T<1$. This gives $\mathcal{K}[u]\in B_{M}$.
Case $p={N(\gamma-1)}/{\theta}$. Fix any $u\in B_{M}$. We take $r>0$ and $k(r)>0$ such that
$1<p/\gamma<r<p=\frac{N(\gamma-1)}{\theta}<k(r)\quad\text{and}\quad\frac{1}{p}=\frac{1/2}{r}+\frac{1/2}{k(r)}.$ (3.3)
$\left(\text{Then inequalities}\quad 1<\frac{q}{\gamma}<\frac{q}{p}r<q<\frac{q}{p}k(r)\quad\text{hold automatically.}\right)$
@@
$\eta=\eta(\lambda):=\lambda^{b},\quad b:=\frac{\theta rk(r)}{N(k(r)-r)}$ (3.4)
@@
$J[u](x,t)$ $=\int_{0}^{\infty}\chi_{\left]0,t\right[}(\tau)S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\,d\tau$
$=\left(\int_{0}^{\eta}+\int_{\eta}^{\infty}\right)\chi_{\left]0,t\right[}(\tau)S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\,d\tau$
$=:J_{\lambda}^{0}[u](x,t)+J_{\lambda}^{1}[u](x,t).$
By using Proposition 2.16, (3.3) and (3.4), we have
$\left\|J_{\lambda}^{0}[u]\mathrel{}\middle|\mathrel{}M^{r}_{qr/p,\infty}\right\|$ $\leqslant\int_{0}^{\eta}\chi_{\left]0,t\right[}(\tau)\left\|S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)\right](x)\mathrel{}\middle|\mathrel{}M^{r}_{qr/p,\infty}\right\|\,d\tau$
$\leqslant C\int_{0}^{\eta}\left(1+\tau^{\frac{N}{\theta}\left(\frac{1}{r}-\frac{\gamma}{p}\right)}\right)\chi_{\left]0,t\right[}(\tau)\left\||u|^{\gamma-1}u(\cdot,t-\tau)\mathrel{}\middle|\mathrel{}M^{p/\gamma}_{q/\gamma,\infty}\right\|d\tau$
$=C\eta^{\frac{N}{\theta}\left(\frac{1}{r}-\frac{\gamma}{p}\right)+1}\sup_{0<t<T}\left\|u(\cdot,t)\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|^{\gamma}\leqslant C\lambda^{1/2}M^{\gamma},$
and similarly we have
$\left\|J_{\lambda}^{1}[u]\mathrel{}\middle|\mathrel{}M^{k(r)}_{{qk(r)}/{p},\infty}\right\|$ $\leqslant C\int_{\eta}^{\infty}\left(1+\tau\right)^{\frac{N}{\theta}\left(\frac{1}{k(r)}-\frac{\gamma}{p}\right)}\chi_{\left]0,t\right[}(\tau)\left\||u|^{\gamma}(\cdot,t-\tau)\mathrel{}\middle|\mathrel{}M^{p/\gamma}_{q/\gamma,\infty}\right\|\,d\tau$
$=C\eta^{\frac{N}{\theta}\left(\frac{1}{k(r)}-\frac{\gamma}{p}\right)+1}\sup_{0<t<T}\left\|u(\cdot,t)\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|^{\gamma}\leqslant C\lambda^{-1/2}M^{\gamma}.$
Proposition 2.8 in the form $M^{p}_{q,\infty}\supset\left(M^{r}_{qr/p,\infty},M^{k(r)}_{qk(r)/p,\infty}\right)_{1/2,\infty}$ implies that
$\left\|J[u](\cdot,t)\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ $\leqslant C\sup_{\lambda>0}\lambda^{-1/2}\inf_{J[\mu]=J_{0}+J_{1}}\left\{\left\|J_{0}\mathrel{}\middle|\mathrel{}M^{r}_{qr/p,\infty}\right\|+\lambda\left\|J_{1}\mathrel{}\middle|\mathrel{}M^{k(r)}_{qk(r)/p,\infty}\right\|\right\}$
$\leqslant C\sup_{\lambda>0}\left\{\lambda^{-1/2}\left\|J_{\lambda}^{0}[u]\mathrel{}\middle|\mathrel{}M^{r}_{qr/p,\infty}\right\|+\lambda^{1/2}\left\|J_{\lambda}^{1}[u]\mathrel{}\middle|\mathrel{}M^{k(r)}_{qk(r)/p,\infty}\right\|\right\}$
@@
$\sup_{0<t<T}\left\|J[u](\cdot,t)\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant\frac{M}{2}$
@@
Fix $0<T<1$. Assume that $\mu$ and $(p,q)$ satisfy the same conditions as in Lemma 3.3. Then the map $\mathcal{K}:B_{M}\to B_{M}$, $u\mapsto\mathcal{K}[u;\mu]$ is a contraction map.
@@
We start by preparing the following estimate, in view of Lemma 2.2. For all $\rho\geqslant\sigma>1$, $\nu>1$ and for all $f$, $g\in M^{\rho\nu}_{\sigma\nu,\infty}$, we have
$\left\||f|^{\nu-1}f-|g|^{\nu-1}g\mathrel{}\middle|\mathrel{}M^{\rho}_{\sigma,\infty}\right\|$
$\qquad\qquad\leqslant\nu\left\|\max\{|f|;|g|\}^{\nu-1}|f-g|\mathrel{}\middle|\mathrel{}M^{\rho}_{\sigma,\infty}\right\|$
$\qquad\qquad\leqslant\nu\left\|\max\{|f|;|g|\}^{\nu-1}\mathrel{}\middle|\mathrel{}M^{\frac{\nu\rho}{\nu-1}}_{\frac{\nu\sigma}{\nu-1},\infty}\right\|\left\|f-g\mathrel{}\middle|\mathrel{}M^{\nu\rho}_{\nu\sigma,\infty}\right\|$
$\qquad\qquad\leqslant\nu\left(\left\|f\mathrel{}\middle|\mathrel{}M^{\nu\rho}_{\nu\sigma,\infty}\right\|^{\nu-1}+\left\|g\mathrel{}\middle|\mathrel{}M^{\nu\rho}_{\nu\sigma,\infty}\right\|^{\nu-1}\right)\left\|f-g\mathrel{}\middle|\mathrel{}M^{\nu\rho}_{\nu\sigma,\infty}\right\|.$ (3.6)
Case $p=N(\gamma-1)/\theta$. Take arbitrary $u,v\in B_{M}$. Again we take $r$ and $k(r)$ as (3.3) holds. Using Proposition 2.8 in the form $M^{p}_{q,\infty}\supset\left(M^{r}_{qr/p,\infty},M^{k(r)}_{qk(r)/p,\infty}\right)_{1/2,\infty}$, we observe that the estimate
$\left\|\mathcal{K}[u]-\mathcal{K}[v]\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$=C\left\|\left(\int_{0}^{\eta}+\int_{\eta}^{\infty}\right)\chi_{\left]0,t\right[}(\tau)S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)-|v|^{\gamma}v(\cdot,t-\tau)\right]\,d\tau\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$\leqslant C\sup_{\lambda>0}\bigg\{\lambda^{-1/2}\int_{0}^{\eta}\chi_{\left]0,t\right[}(\tau)\left\|S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)-|v|^{\gamma}v(\cdot,t-\tau)\right]\mathrel{}\middle|\mathrel{}M^{r}_{qr/p,\infty}\right\|d\tau$
$\qquad+\lambda^{1/2}\int_{\eta}^{\infty}\chi_{\left]0,t\right[}(\tau)\left\|S(\tau)\left[|u|^{\gamma-1}u(\cdot,t-\tau)-|v|^{\gamma}v(\cdot,t-\tau)\right]\mathrel{}\middle|\mathrel{}M^{k(r)}_{qk(r)/p,\infty}\right\|d\tau\bigg\}$
$\leqslant C\sup_{\lambda>0}\left(\lambda^{-1/2}\int_{0}^{\eta}\tau^{\frac{N}{\theta}\left(\frac{1}{r}-\frac{\gamma}{p}\right)}\,d\tau+\lambda^{1/2}\int_{\eta}^{\infty}\tau^{\frac{N}{\theta}\left(\frac{1}{k(r)}-\frac{\gamma}{p}\right)}\,d\tau\right)$
$\qquad\qquad\qquad\qquad\times\left(\left\|u\mathrel{}\middle|\mathrel{}X_{T}\right\|^{\gamma-1}+\left\|v\mathrel{}\middle|\mathrel{}X_{T}\right\|^{\gamma-1}\right)\left\|u-v\mathrel{}\middle|\mathrel{}X_{T}\right\|$
$\leqslant C\sup_{\lambda>0}\left(\lambda^{-1/2}\eta^{1+\frac{N}{\theta r}-\frac{N\gamma}{\theta p}}+\lambda^{1/2}\eta^{1+\frac{N}{\theta k(r)}-\frac{N\gamma}{\theta p}}\right)M^{\gamma-1}\left\|u-v\mathrel{}\middle|\mathrel{}X_{T}\right\|$
holds for any $\eta>0$, through the use of (3.2) and Proposition 2.16. At this stage, optimizing in $\eta$ and taking sufficiently small $M>0$, we derive
$\left\|\mathcal{K}[u]-\mathcal{K}[v]\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant C\sup_{\lambda>0}(1+1)M^{\gamma-1}\left\|u-v\mathrel{}\middle|\mathrel{}X_{T}\right\|\leqslant\frac{1}{2}\left\|u-v\mathrel{}\middle|\mathrel{}X_{T}\right\|.$ (3.7)
Case $p>N(\gamma-1)/\theta$. Utilizing (3.2) and carrying out a similar procedure as in the proof of Lemma 3.3, the claim follows. ∎
@@
From Lemma 3.3, Lemma 3.4 and the contraction mapping theorem, we only have to show $\left\|I[\mu]\mathrel{}\middle|\mathrel{}X_{T}\right\|\leqslant M/2$. Assertion (1) is proved by Proposition 3.2. We next prove (3). This is established by following calculation. We use Proposition 2.11 and 2.15 to see that
$\left\|I[\mu](\cdot,t)\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ $\leqslant C\int_{0}^{t}\left\|S(t-\tau)\mu\mathrel{}\middle|\mathrel{}B^{0,1}_{p,q,\infty}\right\|\,d\tau$
$\leqslant C\int_{0}^{t}\left\{1+(t-\tau)^{s/\theta}\right\}\left\|\mu\mathrel{}\middle|\mathrel{}B^{s,\infty}_{p,q,\infty}\right\|\,d\tau$
$\leqslant C\left(t+t^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s,\infty}_{p,q,\infty}\right\|\leqslant C\left(T+T^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s,\infty}_{p,q,\infty}\right\|$
@@
$\left\|I[\mu]\mathrel{}\middle|\mathrel{}X_{T}\right\|$ $\leqslant C\left(T+T^{1+s/\theta}\right)\left\|\mu\mathrel{}\middle|\mathrel{}B^{s,\infty}_{p,q,\infty}\right\|\leqslant\frac{M}{2}$ (3.8)
@@
$\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{\frac{Np}{N+p\varepsilon}}_{\frac{Nq}{N+p\varepsilon},\infty}\right\|<\delta,$
where $\delta>0$ is determined later. By definition of $\limsup$, there exists an integer $m\in\mathbb{Z}$ such that $2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{\frac{Np}{N+p\varepsilon}}_{\frac{Nq}{N+p\varepsilon},\infty}\right\|<\delta$ holds for all $j\geqslant m$. Set $\mu_{1}:=\sum_{j=0}^{m}\Delta_{j}\mu$ and $\mu_{2}:=\mu-\mu_{1}$. Then for every $j\in\mathbb{N}$, we have the following identities.
$\Delta_{j}\mu_{1}$ $=\begin{cases}\Delta_{j}\mu&\text{for }j\leqslant m-1,\\ \left(\Delta_{m-1}+\Delta_{m}\right)\Delta_{j}\mu&\text{for }j=m,m+1,\\ 0&\text{for }j\geqslant m+2,\end{cases}$
$\Delta_{j}\mu_{2}$ $=\begin{cases}0&\text{for }j\leqslant m-1,\\ \left(\Delta_{m+1}+\Delta_{m+2}\right)\Delta_{j}\mu&\text{for }j=m,m+1,\\ \Delta_{j}\mu&\text{for }j\geqslant m+2.\end{cases}$
Thus we have $\left\|\mu_{2}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta,\infty}_{\frac{Np}{N+p\varepsilon},\frac{Nq}{N+p\varepsilon},\infty}\right\|<C\delta$ and
$\left\|\mu_{1}\mathrel{}\middle|\mathrel{}B^{\sigma,\infty}_{p,q,\infty}\right\|$ $\leqslant C\max_{0\leqslant j\leqslant m+1}2^{\sigma j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant C\max_{0\leqslant j\leqslant m+1}2^{(\sigma+\theta)j}2^{-\theta j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$\leqslant C2^{(\sigma+\theta)(m+1)}\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta,\infty}_{p,q,\infty}\right\|\leqslant C2^{(\sigma+\theta)(m+1)}\left\|\mu\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta,\infty}_{\frac{Np}{N+p\varepsilon},\frac{Nq}{N+p\varepsilon},\infty}\right\|$
for some $\sigma\in\left]-\theta,0\right[$. Now applying (3.8) and Proposition 3.1 with $r=Np/(N+p\varepsilon)$ to $\mu_{1}$ and $\mu_{2}$ respectively, we get
@@
$\leqslant C\left(T+T^{1+\sigma/\theta}\right)\left\|\mu_{1}\mathrel{}\middle|\mathrel{}B^{\sigma,\infty}_{p,q,\infty}\right\|+C\left\|\mu_{2}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta,\infty}_{\frac{Np}{N+p\varepsilon},\frac{Nq}{N+p\varepsilon},\infty}\right\|$
$\leqslant C\left(T+T^{1+\sigma/\theta}\right)2^{(\sigma+\theta)m}\left\|\mu\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta,\infty}_{\frac{Np}{N+p\varepsilon},\frac{Nq}{N+p\varepsilon},\infty}\right\|+C\delta.$
@@
Let $p$ and $q$ satisfy the same conditions in Theorem 1.5. If the inhomogeneous term $\mu$ satisfies one of the conditions in Theorem 1.5, we find a number $s<0$ such that $\mu$ belongs to $B^{s,\infty}_{p,q,\infty}\subset B^{s-N/p,\infty}_{\infty}$. Then the solution $u(x,t)$ constructed above satisfy the initial condition in the following sense: $u(x,t)\to 0$ in the weak-$\ast$ topology of the space $B^{s-N/p,\infty}_{\infty}$ as $t\to+0$. Take arbitrary $\psi\in B^{N/p-s,1}_{1}$ and calculate as follows: